OpenStudy (anonymous):

For which operations is the set {0, 1} closed? a. multiplication b. division c. addition d. subtraction

OpenStudy (anonymous):

@dakid88

OpenStudy (misty1212):

HI!!

OpenStudy (anonymous):

@YoungStudier

OpenStudy (misty1212):

do you know what "closed" means?

OpenStudy (anonymous):

yes

OpenStudy (misty1212):

what does it mean?

OpenStudy (anonymous):

it means that any number in the set using one of those operations equals a number in the set

OpenStudy (misty1212):

exactly!

OpenStudy (youngstudier):

So, what do you think the answer is?

OpenStudy (anonymous):

I can choose more than one

OpenStudy (misty1212):

so what we have to do is try it for each

OpenStudy (anonymous):

I know for sure multiplication

OpenStudy (misty1212):

yes for sure

OpenStudy (anonymous):

possibly division

OpenStudy (misty1212):

how about division

OpenStudy (anonymous):

and definitely not addition

OpenStudy (anonymous):

don't know about subtraction

OpenStudy (misty1212):

\[0\div1\\ 0\div 0\\ 1\div 0\\ 1\div1\]

OpenStudy (anonymous):

oh ok

OpenStudy (misty1212):

what are each of those? if they are either 0 or 1, then closed if not, then not

OpenStudy (misty1212):

and for subtraction same idea \[0-0\\ 0-1\\ 1-0\\ 1-1\]check each

OpenStudy (anonymous):

0/1 equals 0

OpenStudy (misty1212):

ok true, what about the rest of those

OpenStudy (anonymous):

0/0 =0 my calculator says that 1 can not divide by 0

OpenStudy (misty1212):

lol glad your calculator says it, because it is true

OpenStudy (misty1212):

you cannot divide by 0

OpenStudy (anonymous):

ok so division is not closed?

OpenStudy (misty1212):

also if your calculator said \(0\div0=0\) that is a MISTAKE

OpenStudy (misty1212):

yes, it is not closed under division

OpenStudy (anonymous):

0-0=0 0-1=-1

OpenStudy (anonymous):

so it is not closed under subtraction either

OpenStudy (misty1212):

yes it is not

OpenStudy (anonymous):

so that leaves me with only that one choice. but I'm always nervous about only choosing one when I can choose all that apply

OpenStudy (misty1212):

not sure what you mean multiplication, yes

OpenStudy (misty1212):

addition, no (since 1+1=2)

OpenStudy (anonymous):

addition, subtraction, and division no

OpenStudy (misty1212):

subtraction no, since 0 - 1 = -1

OpenStudy (misty1212):

yeah right no on those three, yes on multiplication

OpenStudy (anonymous):

so I'm going to go with just a

OpenStudy (anonymous):

I'm going to leave this question open in case I have more questions. It is a five question test but some of them might be really easy

OpenStudy (misty1212):

i would post in a different box

OpenStudy (anonymous):

oh ok